# Guido Franchetti1 and Derek Harland2

### v1

@@ 3 Existence for gauge fixing condition
The Bogomolony equation for $(A,\Phi)$ is equivalent to the anti-self-dual equations for $B$, which are conformally invariant and take the form
$(F^{B})^{0,2}=0=(F^{B})^{2,0},\quad\omega_{K}\wedge F^{B}=0$ (46)
in terms of the complex structure of $H^{2}\times S^{2}$. Their linearisation is
¯$\bar{\partial}^{B}b^{0,1}=0={\partial}^{B}b^{1,0},\quad\omega_{K}\wedge d^{B}b=0.$ (47)
@@
Thus $B$ determines a holomorphic structure on the vector bundle $V$ over $H^{2}\times S^{2}$ and $b^{0,1}$ is a holomorphic section of $\operatorname{End}_{0}(V)\otimes\Lambda^{0,1}$, with the subscript 0 indicating that $b^{0,1}$ is traceless. This section in turn determines a cohomology class in $H^{0,1}(H^{2}\times S^{2},\operatorname{End}_{0}(V))$. In order to solve the gauge fixing condition it will be convenient to choose a gauge in which ${b}^{0,1}$ vanishes at $w=0$. To do so, we first need to determine the cohomology group $H^{0,1}(H^{2}\times S^{2},\operatorname{End}_{0}(V))$.
We can determine the cohomology of $\operatorname{End}_{0}(V)$ using a natural filtration of $V$, i.e. a holomorphic line bundle $L\subset V$. Following [1], this can be constructed either as the unique circle-invariant continuation of the bundle $L^{-}$ over the 2-sphere at $w=0$, or as the space of decaying solutions to a scattering equation along radial lines. We describe both constructions, beginning with the circle action.
@@
The circle action on $H^{2}\times S^{2}$ is $(w,z)\mapsto(\lambda^{2}w,z)$, with $\lambda=e^{it/2}$. The lift of this action to $V$ fixes $B$ and so preserves the holomorphic structure. It therefore extends to a holomorphic action of the monoid $\mathbb{C}^{\ast}_{\leq 1}=\{\lambda\in\mathbb{C}^{\ast}\>:\>|\lambda|\leq 1\}$ (here the constraint on the modulus of $\lambda$ is imposed in order maintain the condition that $|w|<1$). From (10), the action on the subbundles $L^{\pm}$ over the fixed 2-sphere is given by $\lambda^{\mp 2p}$.
@@
A line subbundle of $V$ corresponds to a section of $\mathbb{P}V$. The line subbundle $L^{-}$ over $\{0\}\times S^{2}$ determines a submanifold $C\subset\mathbb{P}V$ whose points are fixed by the action of $\mathbb{C}^{\ast}_{\leq 1}$. The tangent space of $\mathbb{P}V$ at a point in $C$ can be decomposed into weight spaces for the action, and the weights are $-4p$, 00 and $2$. The weight 00 corresponds to the tangent space of $C$, and the weight $-4p$ corresponds to the tangent space of the fibre of $\mathbb{P}V\to H^{2}\times S^{2}$. The stable submanifold of $C$, consisting of points in $\mathbb{P}V$ that converge to $C$ under the action of $\lambda$ as $\lambda\to 0$, has complex dimension 1. This submanifold is a section of $\mathbb{P}V$, because $\mathbb{C}^{\ast}_{\leq 1}$ acts freely and transitively on $H^{2}\times S^{2}$. It hence determines a line subbundle $L\subset V$. The degree of this line bundle $L$ equals the degree of its restriction to $S^{2}$, which is $-n$.
Now we present the alternative scattering description of $L$. From (43) and (44) we see that $s=-\ln|w|$ is an outward-pointing geodesic radial coordinate on $H^{3}$. Then $w=e^{-s+i\theta}$ is holomorphic coordinate on a surface of constant $z$. Holomorphic sections $\sigma$ on this surface satisfy $(\partial_{s}^{B}-i\partial_{\theta}^{B})\sigma=0$. In terms of the monopole, a holomorphic section which is invariant under the circle action corresponds to a solution of Hitchin’s scattering equation $(\frac{d}{ds}+A_{s}-i\Phi)\sigma=0$ along a geodesic ray parametrised by $s>0$. Solutions typically blow up like $e^{ps}$ as $s\to\infty$, but there is a one-dimensional space of solutions that decays like $e^{-ps}$. These solutions asymptote to $L^{-}$, the eigenspace of $\Phi$ with eigenvalue $ip$. The union of all these lines is the bundle $L$, and the Bogomolny equations ensure that it is a holomorphic subbundle.
@@
$F_{1}$ =$=\{\chi\in\mathrm{End}_{0}(V)\>:\>\chi(V)\subset L,\,\chi(L)=0\}$ (49)
$F_{2}$ =$=\{\chi\in\mathrm{End}_{0}(V)\>:\>\chi(L)\subset L\}.$ (50)
Then $F_{i}$ has rank $i$ and $F_{1}$, $F_{2}/F_{1}$ and $F_{3}/F_{2}$ are line bundles of degree $-n$, 00 and $n$. They are isomorphic to the standard bundles $\mathcal{O}(-n)$, $\mathcal{O}$ and $\mathcal{O}(n)$ over $S^{2}$, pulled back to $H^{2}\times S^{2}$. These isomorphisms can be chosen so that $U(1)$ acts as $w\mapsto\lambda^{2}w$ on $H^{2}$ and by multiplication with $\lambda^{4p}$, $1$ and $\lambda^{-4p}$ on the fibres of $\mathcal{O}(-n)$, $\mathcal{O}$ and $\mathcal{O}(n)$ respectively.
@@
The lemma implies that we can find a gauge transformation $\chi\in\Gamma(H^{2}\times S^{2},\operatorname{End}_{0}(V))$ such that ${b}^{0,1}-\bar{\partial}^{B}\chi$ equals a holomorphic function $f$ on $H^{2}$ times a section $\sigma$ of $\Lambda^{0,1}(S^{2})\otimes\mathcal{O}(-n)$. This can be done in such a way that ${b}-d^{B}\chi$ is still $U(1)$-invariant. Recalling that $U(1)$ acts on $\mathcal{O}(-n)$ as multiplication by $\lambda^{4p}$, this means that the function $f$ must satisfy $f(\lambda^{2}w)=\lambda^{4p}f(w)$. The only such holomorphic function is $f(w)=w^{2p}$. So we are able to choose a gauge in which
${b}^{0,1}=w^{2p}\sigma,\quad\sigma\in\Gamma(\Lambda^{0,1}(S^{2})\otimes\operatorname{End}_{0}(V)).$ (57)
Now let us investigate the implications of this choice for the gauge-fixing condition. The operator $\delta^{-}$ of (40) can be rewritten using the second conformal equivalence of (44) as follows:
If ${b}$ satisfies (57) then this simplifies considerably. To see why, note that (58) implies that
The contraction with $dw$ is 00 in the gauge (57) because $b_{\bar{w}}=0$. So the second term is $w^{3}$ times a smooth function. The first term can be simplified using the linearised anti-self-dual equations (47) and the fact that $d\omega_{K}=0$:
Since $2p\geq 1$, the first term in (59) is also $w^{3}$ times a smooth function. Altogether, we have shown that
@@
Suppose that the gauge is chosen so that ${b}$ satisfies (57) in a neighbourhood of $w=0$. Then
@@
Having chosen this initial gauge, we now try to solve the gauge fixing condition 25. In other words, we seek a section $\chi$ of $\operatorname{End}_{0}(V)$ such that
@@
where $\delta^{-}$ is given in (40) and $b$ satisfies (57). We rewrite this using the following result
@@
$-\left(\frac{1+|w|^{2}}{2w}\right)^{2}\triangle_{S}^{B}\chi+2\left(\frac{1+|w|^{2}}{2w}\right)\ast_{S}\left(d\left(\frac{1+|w|^{2}}{2w}\right)\wedge\ast_{S}\,d^{B}\chi\right).$ (66)
@@
$-\left(\frac{1+|w|^{2}}{2w}\right)^{2}\triangle_{S}^{B}\chi+2\left(\frac{1+|w|^{2}}{2w}\right)\ast_{S}\left(d\left(\frac{1+|w|^{2}}{2w}\right)\wedge\ast_{S}\,d^{B}\chi\right)\\ -\left(\frac{1+|w|^{2}}{2w}\right)\chi(\triangle_{S}+2)\left(\frac{1+|w|^{2}}{2w}\right).$ (67)
@@
Lemma 8 shows that $b$ can be chosen so that the right hand side is a smooth function near $S^{2}\subset S^{4}$. Standard theory guarantees that (69) admits a unique smooth solution. In more detail, the operator $\triangle_{S}^{B}+2$ is a self-adjoint second order elliptic operator. Its kernel is 0, because
@@
By the Fredholm alternative, (69) admits a unique solution $\chi$. By uniqueness, this solution is circle-invariant. So $b_{-}=b+d^{B}\chi$ solves the gauge-fixing condition (62) and is circle-invariant. It gives a solution $(a_{-},\phi_{-})$ of the gauge-fixing condition (25) with a minus sign. This proves the existence part of Theorem 3.
Next we investigate the behaviour of the solution near $S^{2}$. From (57) we see that $b^{0,1}$ is equal to $w$ times a smooth section near $w=0$. From (69) we see that $\chi$ is also equal to $w$ times a smooth section near $w=0$. Since $w$ is a holomorphic coordinate, the $(0,1)$ part of $d^{B}\chi$ is also equal to $w$ times a smooth section, and so the $(0,1)$ part of $b_{-}=b+d^{B}\chi$ equals $w$ times a smooth section. This proves part of the following lemma:
@@
The solution $(a_{-},\phi_{-})$ of (25) in Theorems 3 and 4 is such that the (0,1)-part of $b_{-}=a_{-}+\phi_{-}\,d\theta$ is equal to $w$ times a smooth section near $S^{2}\subset S^{4}$.
@@
The statement is already proved in the case of Theorem 3. For Theorem 4 we in addition need to show that the (0,1)-part of $d^{A}\Phi$ is $w$ times a smooth section. The connection $B=A+\Phi\,d\theta$ has curvature $F^{B}=F^{A}+d^{A}\Phi\wedge d\theta$, so
@@
The anti-self-dual equations (46) imply that $F^{B}$ is of type (1,1), so the (0,1)-part of $d^{A}\Phi$ is $-i{w}{\frac{\partial}{\partial{w}}}\lrcorner F^{B}$. Thus $d^{A}\Phi$ is $w$ times a smooth section. ∎
Now we show that the metric integral (23) is finite for the solution $(a_{-},\phi_{-})$ of the gauge fixing condition (25). This integral is proportional to
@@
The integrand is continuous away from $w=0$, so the integral will be finite as long as the integral in a neighbourhood of 00 is finite. The integral on a neighbourhood $0<|w|<\epsilon$ can be written using the Kähler metric (43) as follows:
$-\int_{0<|w|<\epsilon}\operatorname{Tr}(b_{-}\wedge\ast_{H}b_{-})=-2\int_{0<|w|<\epsilon}\left(\frac{1-|w|^{2}}{2|w|}\right)^{2}\operatorname{Tr}(b_{-}^{1,0}\wedge\ast_{K}b_{-}^{0,1}).$ (73)
[…]

### v2

@@ 3.2 Solution on $S^{2}$
To do that, we first notice that $b^{0,1}$ determines a cohomology class in $H^{0,1}(H^{2}\times S^{2},\operatorname{End}_{0}(V))$, with the subscript 0 indicating that $b^{0,1}$ is traceless. In fact, the Bogomolony equation for $(A,\Phi)$ is equivalent to the anti-self-dual equations for $B$, which are conformally invariant and, in terms of the Kähler structure of $H^{2}\times S^{2}$, take the form
$(F^{B})^{0,2}=0=(F^{B})^{2,0},\quad\omega_{K}\wedge F^{B}=0.$ (75)
Their linearisation is
$\bar{\partial}^{B}b^{0,1}=0={\partial}^{B}b^{1,0},\quad\omega_{K}\wedge d^{B}b=0.$ (76)
@@
Thus $B$ determines a holomorphic structure on the vector bundle $V$ over $H^{2}\times S^{2}$ and $b^{0,1}$ is a $\bar{\partial}^{B}$-closed section of $\operatorname{End}_{0}(V)\otimes\Lambda^{0,1}$, which in turn determines a cohomology class in $H^{0,1}(H^{2}\times S^{2},\operatorname{End}_{0}(V))$. Hence in order to find a gauge where (73) is satisfied it is useful to first determine the cohomology group $H^{0,1}(H^{2}\times S^{2},\operatorname{End}_{0}(V))$.
We can determine the cohomology of $\operatorname{End}_{0}(V)$ using a natural filtration of $V$, i.e. a holomorphic line bundle $L\subset V$. Following [1], we show that:
Lemma 18.
There is a unique circle-invariant holomorphic line bundle $L\subset V$ over $H^{2}\times S^{2}$ such that the restriction of $L$ to $\{0\}\times S^{2}$ equals the line bundle $L^{-}$ that appears in the monopole boundary conditions. The degree of this bundle is n-n.
Proof.
@@
The circle action (14) on $H^{2}\times S^{2}$ is $(w,z)\mapsto(\lambda w,z)$, with $\lambda=e^{it}$. This lifts to an action of $S^{1}$ on the bundle $\mathbb{P}V$. The action fixes the holomorphic structure of $\mathbb{P}V$, because the holomorphic structure is determined by the connection $B$ and $B$ is circle-invariant. So the circle action on $\mathbb{P}V$ can be extended to a holomorphic action of the monoid $\mathbb{C}^{\ast}_{\leq 1}=\{\lambda\in\mathbb{C}^{\ast}\>:\>|\lambda|\leq 1\}$ (here the constraint on the modulus of $\lambda$ is imposed in order maintain the condition that $|w|<1$).
@@
A line subbundle of $V$ corresponds to a section of $\mathbb{P}V$. The line subbundle $L^{-}$ determines a section of $\mathbb{P}V$ over $\{0\}\times S^{2}$. Let $C$ be the image of this section. Then $C$ is a submanifold of $\mathbb{P}V$ whose points are fixed by the action of $\mathbb{C}^{\ast}_{\leq 1}$. We will show that the unstable submanifold $C^{u}$ of $C$, which consists of points in $\mathbb{P}V$ that converge to $C$ under the action of $\lambda$ as $\lambda\to 0$, is the image of a section of $\mathbb{P}V$. This section determines a circle-invariant holomorphic line bundle $L\subset V$.
[…]
$0\to\operatorname{Hom}(L^{-}_{(0,z)},L^{+}_{(0,z)})\to T_{c}\mathbb{P}V\to T_{0}H^{2}\oplus T_{z}S^{2}.$ (77)
From proposition (6), the circle group acts on the first space with weight 4p-4p, and it acts on the final two spaces with weights $2$ and 0. So it acts on $T_{c}\mathbb{P}V$ with weights 4p-4p, 0 and $2$.
We see that there are two non-negative weights, 2 and 0. So the unstable manifold $C^{u}$ has complex dimension 2. The tangent space of $C^{u}$ at $c\in C$ is the union of the weight 2 and 0 subspaces of $T_{c}\mathbb{P}V$, and this maps surjectively to $T_{0}H^{2}\oplus T_{z}S^{2}$ in (77).
We need to show that the submanifold $C^{u}\subset\mathbb{P}V$ is the image of a section of $\mathbb{P}V\to H^{2}\times S^{2}$. The inverse function theorem shows that this is true in an neighbourhood $\{w<\epsilon\}$ of $\{0\}\times S^{2}$, because the tangent space at $c\in C$ maps surjectively to the tangent space of $H^{2}\times S^{2}$. So we have a section $\sigma:\{w<\epsilon\}\times S^{2}\to\mathbb{P}(V)$ whose graph is part of $C^{u}$. Using the action of $\lambda=\epsilon$, we can extend this to a section $\sigma(w,z)=\epsilon^{-1}\cdot\sigma_{\epsilon}(\epsilon w,z)$ of $\mathbb{P}V$ on the whole of $H^{2}\times S^{2}$. The image of this section is the submanifold $C^{u}$.
So $C^{u}$ is the graph of a section of $\mathbb{P}V$. It therefore determines a line bundle $L\subset V$. The degree of $L$ equals the degree of its restriction to $S^{2}$, which is n-n. ∎
Remark 19.
The line bundle $L\subset V$ can alternatively be described using monopole scattering [1]. Fix a point $z_{0}\in H^{2}$; this determines a radial line in $H^{3}$. Recall that Hitchin’s scattering equation along this line is
$\left(\frac{d}{ds}+A_{s}-i\Phi\right)\sigma=0,$ (78)
in which $\sigma$ is a section of $V$ over the line and $s=-\ln|w|$ is the geodesic radial coordinate on $H^{3}$. Then the fibre of $L$ above $(w_{0},z_{0})\in H^{2}\setminus\{0\}\times S^{2}$ is the set of $\sigma_{0}\in V_{w_{0},z_{0}}$ such that the solution of the scattering equation with initial condition $\sigma=\sigma_{0}$ at $s=s_{0}:=-\ln|w_{0}|$ decays as $s\to\infty$.
To explain why this characterisation of $L$ is equivalent to that of Lemma 18, we note that if $\sigma$ is a solution of (78) and satisfies $\partial_{\theta}\sigma=0$, then it satisfies $(\partial_{s}^{B}-i\partial_{\theta}^{B})\sigma=0$ for the connection $B=\Phi\,d\theta+A$. Since $w=e^{-s+i\theta}$, this means that $\sigma$ determines a circle-invariant holomorphic section of $V$ over $H^{2}\setminus\{0\}\times\{z_{0}\}$. If $\sigma$ decays as $s\to\infty$ then it must approach the eigenspace of $\Phi|_{S^{2}_{\infty}}$ with eigenvalue $ip$. But this is precisely the fibre of $L^{-}$ over $z_{0}\in S^{2}_{\infty}$. So $\sigma$ must be a section of the circle-invariant bundle $L|_{H^{2}\setminus\{0\}\times\{z_{0}\}}$ that extends $L^{-}_{z_{0}}$. ∎
@@
$F_{1}$ $=\{\chi\in\mathrm{End}_{0}(V)\>:\>\chi(V)\subset L,\,\chi(L)=0\}$ (80)
$F_{2}$ $=\{\chi\in\mathrm{End}_{0}(V)\>:\>\chi(L)\subset L\}.$ (81)
Then $F_{i}$ has rank $i$ and $F_{1}$, $F_{2}/F_{1}$ and $F_{3}/F_{2}$ are line bundles of degree n-n, 0 and $n$. They are isomorphic to the standard bundles $\mathcal{O}(-n)$, $\mathcal{O}$ and $\mathcal{O}(n)$ over $S^{2}$, pulled back to $H^{2}\times S^{2}$. These isomorphisms can be chosen so that $U(1)$ acts as $w\mapsto\lambda w$ on $H^{2}$ and by multiplication with $\lambda^{2p}$, $1$ and $\lambda^{-2p}$ on the fibres of $\mathcal{O}(-n)$, $\mathcal{O}$ and $\mathcal{O}(n)$ respectively.
@@
Let now $\sigma_{1},\sigma_{2},\ldots$ be sections of $\Lambda^{0,1}(S^{2})\otimes\mathcal{O}(-n)$ that represent a basis for $H^{0,1}(S^{2},\mathcal{O}(-n))$. By the preceding lemma we can find an infinitesimal gauge transformation $\chi\in\Gamma(H^{2}\times S^{2},\operatorname{End}_{0}(V))$ and holomorphic functions $f_{i}:H^{2}\to\mathbb{C}$ such that
${b}^{0,1}-\bar{\partial}^{B}\chi=\sum_{i}f_{i}\sigma_{i}.$ (88)
This can be done in such a way that ${b}-d^{B}\chi$ is still $U(1)$-invariant. Recalling that $U(1)$ acts on $\mathcal{O}(-n)$ with weight $2p$, this means that
$\sum_{i}f_{i}\sigma_{i}=\sum_{i}\lambda^{2p}f_{i}(\lambda^{-1}w)\sigma_{i}(u)\quad\forall\lambda\in U(1).$ (89)
Since the $\sigma_{i}$ are linearly independent, it follows that $\lambda^{2p}f_{i}(\lambda^{-1}w)=f_{i}(w)$ for each $i$. The only holomorphic functions with this property are constant multiples of $w\mapsto w^{2p}$. So the right hand side of (88) is $w^{2p}\sigma$ for some linear combination $\sigma$ of the $\sigma_{i}$. Therefore we are able to make an infinitesimal gauge transformation so that
$b^{0,1}=w^{2p}\sigma,\quad\sigma\in\Gamma(\Lambda^{0,1}(S^{2})\otimes\operatorname{End}_{0}(V)),$ (90)
where, abusing notation, we have denoted by $b$ the gauge transformed deformation previously denoted by $b-d^{B}\chi$.
With this choice of gauge, $\delta^{-}b$ decays as $w\to 0$:
@@
Suppose that the gauge is chosen so that b{b} satisfies (90) in a neighbourhood of $w=0$. Then
@@
Proof.
The operator $\delta^{-}$ of (65) can be rewritten using the second conformal equivalence of (71) as follows:
If b{b} satisfies (90) then this simplifies considerably. To see why, note that (92) implies that
The contraction with $dw$ is 0 in the gauge (90) because $b_{\bar{w}}=0$ and $dw\lrcorner dw=0$. So the second term is $w^{3}$ times a smooth function. The first term can be simplified using the linearised anti-self-dual equations (76) and the fact that $d\omega_{K}=0$:
Since $2p\geq 1$, the first term in (93) is also $w^{3}$ times a smooth function. ∎
3.3 Solution on $S^{4}$
Having found a gauge where (90) holds, we now try to solve the gauge fixing condition (64) on $S^{4}$. In other words, we seek a section $\chi$ of $\operatorname{End}_{0}(V)\to S^{4}$ such that
@@ 3.3 Solution on $S^{4}$
where $\delta^{-}$ is given in (65). We rewrite (95) using the following result
@@
$\left(\frac{1+|w|^{2}}{2w}\right)^{2}\triangle_{S}^{B}\chi+2\left(\frac{1+|w|^{2}}{2w}\right)\ast_{S}\left(d\left(\frac{1+|w|^{2}}{2w}\right)\wedge\ast_{S}\,d^{B}\chi\right).$ (99)
@@
$\left(\frac{1+|w|^{2}}{2w}\right)^{2}\triangle_{S}^{B}\chi+2\left(\frac{1+|w|^{2}}{2w}\right)\ast_{S}\left(d\left(\frac{1+|w|^{2}}{2w}\right)\wedge\ast_{S}\,d^{B}\chi\right)\\ -\left(\frac{1+|w|^{2}}{2w}\right)\chi(\triangle_{S}+2)\left(\frac{1+|w|^{2}}{2w}\right).$ (100)
@@
Lemma 21 shows that, since $b$ satisfies (90), the right hand side is a smooth function near $S^{2}\subset S^{4}$. Standard theory then guarantees that (102) admits a unique smooth solution. In more detail, the operator $\triangle_{S}^{B}+2$ is a self-adjoint second order elliptic operator. Its kernel is 0, because
@@
By the Fredholm alternative, (102) admits a unique solution $\chi$. By uniqueness, this solution is circle-invariant. So $b_{-}=b+d^{B}\chi$ solves the gauge-fixing condition (95) and is circle-invariant. It gives a solution $(a_{-},\phi_{-})$ of the gauge-fixing condition (49) with a minus sign.
To complete the proof of the existence part of Theorems 12, 13 we still need to show that $a_{-}$ is a section of $\Lambda^{1,0}$ over $S^{2}_{\infty}$ and that the $L^{2}$ norm of $(a_{-},\phi_{-})$ is finite. Let us start with the behaviour over $S^{2}_{\infty}$.
@@
The solution $(a_{-},\phi_{-})$ of (49) in Theorems 12 and 13 is such that the (0,1)-part of $b_{-}=a_{-}+\phi_{-}\,d\theta$ is equal to $w$ times a smooth section near $S^{2}\subset S^{4}$.
@@
From (90) we see that $b^{0,1}$ is equal to $w$ times a smooth section near $w=0$. From (102) we see that $\chi$ is also equal to $w$ times a smooth section near $w=0$. Since $w$ is a holomorphic coordinate, the $(0,1)$ part of $d^{B}\chi$ is also equal to $w$ times a smooth section, and so the $(0,1)$ part of $b_{-}=b+d^{B}\chi$ equals $w$ times a smooth section. This proves the Lemma in the case of Theorem 12. For Theorem 13 we need in addition to show that the (0,1)-part of $d^{A}\Phi$ is $w$ times a smooth section. The connection $B=A+\Phi\,d\theta$ has curvature $F^{B}=F^{A}+d^{A}\Phi\wedge d\theta$, so
@@
The anti-self-dual equations (75) imply that $F^{B}$ is of type (1,1), so the (0,1)-part of $d^{A}\Phi$ is $i{w}{\frac{\partial}{\partial{w}}}\lrcorner F^{B}$. Thus $d^{A}\Phi$ is $w$ times a smooth section. ∎
Finally, we show that the solution $(a_{-},\phi_{-})$ of the gauge fixing condition (49) has finite $L^{2}$ norm. The $L^{2}$ integral is proportional to
@@
The integrand is continuous away from $w=0$, so the integral will be finite as long as the integral in a neighbourhood of 0 is finite. The integral on a neighbourhood $0<|w|<\epsilon$ can be written using the Kähler metric (69) as follows:
$\int_{0<|w|<\epsilon}\operatorname{Tr}(b_{-}\wedge\ast_{H}b_{-})=-2\int_{0<|w|<\epsilon}\left(\frac{1-|w|^{2}}{2|w|}\right)^{2}\operatorname{Tr}(b_{-}^{1,0}\wedge\ast_{K}b_{-}^{0,1}).$ (106)
By Lemma 23, $b_{-}^{0,1}$ is $w$ times a smooth section, so the integrand is $dw\wedge d\bar{w}/\bar{w}$ times the volume form of $S^{2}$ times a function that extends smoothly over $w=0$. Since $\int_{0<|w|<\epsilon}dw\wedge d\bar{w}/|w|$ converges, the integral (106) converges. This completes the proof of Theorem 12.